OpenStudy (anonymous):

x^(1-a) / x^a = ? Show work? I really don't know how to do this onee.

OpenStudy (amistre64):

\[x^{1-a} = x^1 * x^{-a}\]

OpenStudy (anonymous):

1/x^a = x^-a x^(1-a)x^-a When you multiply terms, you add the exponents. x^(1-2a)

OpenStudy (anonymous):

\[\frac{x^{1-a}}{x^a}=x^{1-a-a}=x^{1-2a}\]